• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 17:29:46 on September17,2009

    Entry number 290163

    keyword=beam checked, production resumed with 98 uA.

    After backing out of Moller mode, MCC got us beam back, did a harp scan (both 1H03A&B) as we requested. They look normal. We then got 40 uA on empty target did spot++ with raser off first, then raster on
    (got 6x4.5mm). Moved to LH2 target, requested 98 uA, resumed production run. Checked Lumi width, they are reasonable (1145-1200).
    Compton is also running now.